# Number it: Temporal Grounding Videos like Flipping Manga

### v2

@@ Abstract
Video Large Language Models (Vid-LLMs) have made remarkable advancements in comprehending video content for QA dialogue. However, they struggle to extend this visual understanding to tasks requiring precise temporal localization, known as Video Temporal Grounding (VTG). To address this gap, we introduce Number-Prompt (NumPro), a novel method that empowers Vid-LLMs to bridge visual comprehension with temporal grounding by adding unique numerical identifiers to each video frame. Treating a video as a sequence of numbered frame images, NumPro transforms VTG into an intuitive process: flipping through manga panels in sequence. This allows Vid-LLMs to “read” event timelines, accurately linking visual content with corresponding temporal information. Our experiments demonstrate that NumPro significantly boosts VTG performance of top-tier Vid-LLMs without additional computational cost. Furthermore, fine-tuning on a NumPro-enhanced dataset defines a new state-of-the-art for VTG, surpassing previous top-performing methods by up to 6.9% in mIoU for moment retrieval and 8.5% in mAP for highlight detection. The code will be available at https://github.com/yongliang-wu/NumPro.
@@ 1 Introduction
Imagine you are watching a cooking video, and trying to locate the exact moment when the chef stirs in the spices. While recognizing such actions is feasible, translating that visual information into precise timing, i.e., a specific second or frame number, is surprisingly difficult. This challenge is central to the field of Video Temporal Grounding (VTG) [17, 4, 43, 47, 22]. In the realm of Video Large Language Models (Vid-LLMs) [55, 36, 71, 29], the integration of VTG allows for fine-grained visual and temporal understanding and reasoning of videos, which is pivotal for developing end-to-end video dialogue systems.
Despite advances of Vid-LLMs, endowing these models with effective VTG abilities presents a unique challenge: enhancing the model’s visual recognition of an event within a video does not inherently enable it to describe when the event begins and ends using language [22, 47]. For instance, advanced Vid-LLMs like Qwen2-VL [55], while excelling at video comprehension, can struggle with grounding specific events in time. When asked, e.g., to locate “when does the woman eat food” in a 10-frame video, the model can hallucinate an illogical answer like “from frame 000 to 580.” [^1] See more cases in Appendix 7.1. This limitation arises because these models are primarily trained to align visual content with language descriptions (what happens) while lacking mechanisms to directly interpret the temporal boundaries (when does it happen).
@@
Current Vid-LLMs process videos by considering them as a sequence of frame images. With NumPro, VTG is as intuitive as flipping manga. As shown in Figure 1, NumPro augments each frame with a unique numerical identifier denoting its position in the temporal sequence. Given a language query targeting an event, Vid-LLMs retrieve relevant visual features of video frames and associate with the frame numbers overlaid. These numerical identifiers are then directly translated into textual outputs. In practice, we strategically position frame numbers in the bottom-right corner, using a medium font size and distinct color. This design ensures numbers visibility without obstructing essential visual content. Overall, NumPro allows Vid-LLMs to “read” the video timeline, effectively converting visual recognition into a temporal narrative in language.
@@
  • We introduce NumPro, a novel approach that enhances Video Temporal Grounding (VTG) capabilities of Video Large Language Models (Vid-LLMs) by overlaying frame numbers onto video frames, making temporal grounding as intuitive as following numbered panels in flipping manga.
@@
2 Related Work
Video Temporal Grounding with Vid-LLMs. Video Temporal Grounding (VTG) [43, 54, 72, 32] focuses on the precise identification of event timestamps within videos, covering tasks such as moment retrieval [17, 16, 3, 68, 53, 4, 7, 63, 14, 62, 37], dense captioning [26, 24, 64, 53, 4, 8], and highlight detection [28, 47]. For current Video Large Language Models (Vid-LLMs) [55, 71, 29], which leverage powerful LLMs [1] for cross-modal understanding and video-based reasoning, VTG is crucial for achieving fine-grained temporal and visual comprehension, enabling end-to-end video dialogue systems with integrated temporal reasoning [22, 47, 56]. To achieve this, some methods rely on refined instruction datasets with temporal information (timestamps or frame numbers) for model fine-tuning [22, 33], while others concatenate additional textual temporal timestamps tokens with visual inputs [20, 50] or introduce specific temporal embeddings [47, 18]. Additional strategies model video structure [19, 23, 56] to better segment or organize videos into parts suitable for VTG. However, these approaches often require extensive retraining or specialized model adaptations, limiting their flexibility and transferability. In contrast, our NumPro aims to improve VTG for existing Vid-LLMs without additional training costs or architectural modifications.
@@ 3.1 Attention Analysis
Specifically, we take a video and a language query as input, and extract the attention scores from the final multi-head self-attention layer of Qwen2-VL-7B [55]. For each frame within the video sequence, we aggregate the attention scores from all the visual tokens corresponding to that frame across all attention heads.
As illustrated in Figure 2, the attention map reveals a strong correlation between the text query of an event and targeted video segments. It indicates that Qwen2-VL-7B can effectively focus on query-relevant frames, which is consistent with the model’s strong performance in other content-related video understanding tasks [15, 30]. However, the model struggles to verbalize the correct temporal boundaries, and generates surprising hallucinations such as “from 200 to 599.” [^2] See more cases in Appendix 8. This observation underscores the need for mechanisms that can bridge the gap between spatial feature alignment and temporal reasoning with Vid-LLMs, which we aim to address with our NumPro method.
@@ 3.2 NumPro and NumPro-FT
In the training-free setting, each video frame is marked with its corresponding frame number. By utilizing the built-in Optical Character Recognition (OCR) capabilities of Vid-LLMs, we enable them to “read” the timeline through the frame numbers associated with visual content. To clarify the purpose of the added numbers to Vid-LLMs, we prepend a simple instruction to each event query: “The red numbers on each frame represent the frame number.” This approach allows Vid-LLMs to accurately identify frame-level boundaries by directly linking the frame numbers to language queries.
@@ 3.3 Design of Numerical Prompt
To make the design choices robust across various models and datasets, we employ CLIP-based experiments on a subset of MSCOCO dataset [35] to calculate Number Accuracy and Caption Accuracy separately. We use the CLIP ViT-B/32 [46, 11] model to generate visual and textual representations, as many Vid-LLMs utilize CLIP-style vision encoders [55, 12], allowing our findings to generalize well across Vid-LLMs. COCO image-caption pairs serve as proxies for video frames, avoiding the high costs and limited scalability of direct VTG testing. Specifically, we randomly select 1,000 distinct image-caption pairs from MSCOCO [35] and overlay numbers ranging from “0” to “99” onto the image in various configurations.
@@
Figure 4: Illustration of Our NumPro Design Algorithm.We overlay different numbers onto COCO images and obtain visual and textual representations using CLIP encoders. For each configuration, we calculate Number/Caption Similarity and derive Number/Caption Accuracy, enabling us to identify the optimal NumPro design that balances recognizability and minimal disruption to the visual content.
@@
As shown in Figure 5, our findings indicate that increasing the font size improves number accuracy but reduces caption accuracy, suggesting that a moderate font size (40 or 60) is optimal. For color selection, caption accuracy remains relatively stable across different colors. Red shows the best performance for number accuracy, while black was the least effective. This finding is also consistent with previous works [49, 5]. Additionally, positioning the text in the center of the image significantly reduced caption accuracy due to overlaps with key visual elements, while placing the numbers in the bottom-right corner provides the best balance between caption and number accuracy. Finally, we select a font size of 40, the color red, and the bottom-right position for our final NumPro design. This design search allows NumPro to better harness the inherent OCR and visual-language alignment capabilities of Vid-LLMs to empower video temporal grounding.
@@ 4.1 Implementation Details for NumPro-FT
Training Details. We utilize the LongVA-7B-DPO [69] as our base model, taking into account its uncomplicated design and its extensive capacity to handle context length. Additionally, it has not been trained on any video data. The model is trained for 3 epochs over our curated dataset with a total batch size of 128. We use the AdamW optimizer [25] with cosine learning rate decay. The learning rate is set to 1e-4, and the warm-up ratio is 0.05. The LLM component utilizes LoRA with parameters $r=64$ and $\alpha=128$. All experiments are conducted on 8 H100 GPUs.
@@ 4.3 Validation of NumPro Design
Following our heuristic design process in Sec 3.3, we validate its effectiveness in temporal grounding tasks to confirm that these design choices generalize beyond the COCO dataset. We conduct moment retrieval experiments on Charades-STA [17] with Qwen2-VL-7B [55] in a training-free setting. As shown in Table 3, the results align closely with our initial observations from COCO dataset, confirming the effectiveness of our design choices in VTG tasks. Specifically, (1) Position: Consistent with our CLIP-based findings, placing the text in the center has the largest impact on performance due to overlaps, while our choice of the bottom-right performs comparably to the best position; (2) Font Size: Both very large and very small fonts yield suboptimal results, supporting our balanced selection; (3) The performance on VTG is sensitive to number color, yet the red color consistently delivers the best performance, which may attribute to its high contrast against typical backgrounds [49]. Overall, the alignment between the CLIP-based design choices and the VTG results shows the validity and robustness of our NumPro design. Please refer to Appendix 10 for the ablation results of NumPro-FT.
@@ 4.5 Influence on General Video-QA
To explore the broader applicability of NumPro, we integrate it into general video-QA tasks, using VideoInstruct [41] as our benchmark. As detailed in Table 4, the incorporation of NumPro minimally affects general comprehension metrics, with a slight decrease in Distraction Overlap (DO, -0.02) and an enhancement in Temporal Understanding (TU, +0.1). This indicates that Vid-LLMs equipped with NumPro maintain robust performance in general video-QA while excelling in precise video temporal grounding (VTG) tasks. This dual capability allows us to harness a powerful Vid-LLM for end-to-end video understanding that can flexibly adapt to both general and temporally nuanced questions within conversational AI systems.
@@ 7.2 VTG-Tuned Vid-LLMs
Table 5: The ablation results on QVHighlights dataset.

### v3

@@ Abstract
Video Large Language Models (Vid-LLMs) have made remarkable advancements in comprehending video content for QA dialogue. However, they struggle to extend this visual understanding to tasks requiring precise temporal localization, known as Video Temporal Grounding (VTG). To address this, we introduce Number-Prompt (NumPro), a novel method that empowers Vid-LLMs to bridge visual comprehension with temporal grounding by adding unique numerical identifiers to each video frame. Treating a video as a sequence of numbered frame images, NumPro transforms VTG into an intuitive process: flipping through manga panels in sequence. This allows Vid-LLMs to “read” event timelines, accurately linking visual content with corresponding temporal information. Our experiments demonstrate that NumPro significantly boosts VTG performance of top-tier Vid-LLMs without additional computational cost. Furthermore, fine-tuning on a NumPro-enhanced dataset defines a new state-of-the-art for VTG, surpassing previous top-performing methods by up to 6.9% in mIoU for moment retrieval and 8.5% in mAP for highlight detection. The code is available at https://github.com/yongliang-wu/NumPro.
@@
footnotetext: 11footnotemark: 1 Equal Contribution.  33footnotemark: 3 Project Leader.  44footnotemark: 4 Corresponding Author.
22footnotemark: 2 Work done during an internship at WeChat Vision, Tencent Inc.
@@ 1 Introduction
Imagine you are watching a cooking video, and trying to locate the exact moment when the chef stirs in the spices. While recognizing such actions is feasible, translating that visual information into precise timing, i.e., a specific second or frame number, is surprisingly difficult. This challenge is central to the field of Video Temporal Grounding (VTG) [18, 4, 52, 58, 25, 36]. In the realm of Video Large Language Models (Vid-LLMs) [54, 66, 43, 89, 35, 84] which process videos as a sequence of frame images, the integration of VTG allows for fine-grained visual and temporal understanding and reasoning of videos, which is pivotal for developing end-to-end video dialogue systems.
Despite advances of Vid-LLMs, endowing these models with effective VTG abilities presents a unique challenge: enhancing the model’s visual recognition of an event within a video does not inherently enable it to describe when the event begins and ends using language [25, 58]. For instance, advanced Vid-LLMs like Qwen2-VL [66], while excelling at video comprehension, can struggle with grounding specific events in time. When asked, e.g., to locate “when does the woman eat food” in a 10-frame video, the model can hallucinate an illogical answer like “from frame 000 to 580.”***See more cases in Appendix 7.1. This limitation arises because these models are primarily trained to align visual content with language descriptions (what happens) while lacking mechanisms to directly interpret the temporal boundaries (when does it happen). å
@@
With NumPro, VTG is as intuitive as flipping manga. As shown in Figure 1, NumPro augments each frame with a unique numerical identifier denoting its position in the temporal sequence. Given a language query targeting an event, Vid-LLMs retrieve relevant visual features of video frames and associate them with the frame numbers overlaid. These numerical identifiers are then directly translated into textual outputs. In practice, we strategically position frame numbers in the bottom-right corner, using a defined font size and distinct color. This design ensures numbers visibility without obstructing essential visual content. Overall, NumPro allows Vid-LLMs to “read” the video timeline, effectively converting visual recognition into a temporal narrative.
@@
  • We introduce NumPro, a novel approach that enhances Video Temporal Grounding (VTG) capabilities of Vid-LLMs by overlaying frame numbers onto video frames, making temporal grounding as intuitive as following numbered panels in flipping manga.
@@
2 Related Work
Video Temporal Grounding with Vid-LLMs. Video Temporal Grounding (VTG) [52, 65, 90, 39] focuses on the precise identification of event timestamps within videos, covering tasks such as moment retrieval [18, 17, 3, 86, 64, 4, 7, 76, 15, 75, 45, 83, 81, 82, 34], dense captioning [31, 28, 77, 64, 4, 9, 19, 85, 22, 56, 67], and highlight detection [33, 58]. For current Video Large Language Models (Vid-LLMs) [66, 89, 35], which leverage powerful LLMs [1] for cross-modal understanding and video-based reasoning, VTG is crucial for achieving fine-grained temporal and visual comprehension, enabling end-to-end video dialogue systems with integrated temporal reasoning [25, 58, 68]. To achieve this, some methods rely on refined instruction datasets with temporal information (timestamps or frame numbers) for model fine-tuning [25, 40], while others concatenate additional textual temporal timestamps tokens with visual inputs [23, 61] or introduce specific temporal embeddings [58, 20]. Additional strategies model video structure [21, 26, 68] to better segment or organize videos into parts suitable for VTG. However, these approaches often require extensive retraining or specialized model adaptations, limiting their flexibility and transferability. In contrast, our NumPro aims to improve VTG for existing Vid-LLMs without additional training costs or architectural modifications.
@@ 3.1 Attention Analysis
Specifically, we take a video and a language query as input, and extract the attention scores from the final multi-head self-attention layer of Qwen2-VL-7B [66]. For each frame within the video sequence, we aggregate the attention scores from all the visual tokens corresponding to that frame across all attention heads. As illustrated in Figure 2, the attention map reveals a strong correlation between the text query of an event and targeted video segments. It indicates that Qwen2-VL-7B can effectively focus on query-relevant frames, which is consistent with the model’s strong performance in other content-related video understanding tasks [16, 37, 74]. However, the model struggles to verbalize the correct temporal boundaries, and generates surprising hallucinations such as “from 200 to 599.”See more cases in Appendix 8. This observation underscores the need for mechanisms that can bridge the gap between spatial feature alignment and temporal reasoning with Vid-LLMs, which we aim to address.
@@ 3.2 NumPro and NumPro-FT
In the training-free setting, each video frame is marked with its corresponding frame number. By utilizing the built-in Optical Character Recognition (OCR) capabilities of Vid-LLMs, we enable them to “read” the timeline through the frame numbers associated with visual content. To clarify the purpose of the added numbers to Vid-LLMs, we prepend a simple instruction to each event query: “The red numbers on each frame represent the frame number.” This approach allows Vid-LLMs to identify frame-level boundaries by directly linking the frame numbers to language queries.
@@ 3.3 Design of Numerical Prompt
To make the design choices robust across various models and datasets, we employ CLIP-based experiments on a subset of MSCOCO dataset [42] to calculate Number Accuracy and Caption Accuracy separately. We use the CLIP ViT-B/32 [57, 12, 46, 8, 91, 30, 27] model to generate visual and textual representations, as many Vid-LLMs utilize CLIP-style vision encoders [66, 13, 44], allowing our findings to generalize well across Vid-LLMs. COCO image-caption pairs serve as proxies for video frames, avoiding the high costs and limited scalability of direct VTG testing. Specifically, we randomly select 1,000 distinct image-caption pairs from MSCOCO [42] and overlay numbers ranging from “0” to “99” onto the image in various configurations.
@@
Figure 4: Our NumPro Design Algorithm. We overlay different numbers onto COCO images and obtain visual and textual representations using CLIP encoders. For each configuration, we calculate Number/Caption Similarity and derive Number/Caption Accuracy, to identify the optimal NumPro design that balances recognizability and minimal disruption to the visual content.
@@
As shown in Figure 5, our findings indicate that increasing the font size improves number accuracy but reduces caption accuracy, suggesting that a moderate font size (40 or 60) is optimal. For color selection, caption accuracy remains relatively stable across different colors. Red shows the best performance for number accuracy, while black was the least effective. This finding is also consistent with previous works [60, 5]. Additionally, positioning the text in the center of the image significantly reduced caption accuracy due to overlaps with key visual elements, while placing the numbers in the bottom-right corner provides the best balance between caption and number accuracy. Finally, we select a font size of 40, the color red, and the bottom-right position for our final NumPro design.
@@ 4.1 Implementation Details for NumPro-FT
Training Details. We utilize the LongVA-7B-DPO [87] as our base model, taking into account its uncomplicated design and its extensive capacity to handle context length. Additionally, it has not been trained on any video data. The model is trained for 3 epochs over our curated dataset with a total batch size of 128. We use the AdamW optimizer [29] with cosine learning rate decay. The learning rate is set to 1e-4, and the warm-up ratio is 0.05. The LLM component utilizes LoRA with parameters $r=64$ and $\alpha=128$. All experiments are conducted on 8 H800 GPUs.
@@ 4.3 Validation of NumPro Design
Following our heuristic design process in Sec 3.3, we validate its effectiveness in temporal grounding tasks to confirm that these design choices generalize beyond the COCO dataset. We conduct moment retrieval experiments on Charades-STA [18] with Qwen2-VL-7B [66] in a training-free setting. As shown in Table 3, the results align closely with our initial observations from the COCO dataset, confirming the effectiveness of our design choices in VTG tasks. Specifically, (1) Position: Consistent with our CLIP-based findings, placing the text in the center has the largest impact on performance due to overlaps, while our choice of the bottom-right performs comparably to the best position; (2) Font Size: Both very large and very small fonts yield suboptimal results, supporting our balanced selection; (3) The performance on VTG is sensitive to number color, yet the red color consistently delivers the best performance, which may attribute to its high contrast against typical backgrounds [60]. Overall, the alignment between the CLIP-based design choices and the VTG results shows the validity and robustness of our NumPro design. Please refer to Appendix 11 for the ablation results of NumPro-FT. We also try directly overlaying timestamps (e.g., “10.5s”) on frames, which show inferior performance than frame numbers (Appendix 12).
@@ 4.5 Influence on General Video-QA
To explore the broader applicability of NumPro, we integrate it into general video-QA tasks, using VideoInstruct [50] as our benchmark. As detailed in Table 4, the incorporation of NumPro minimally affects general comprehension metrics, with a slight decrease in Distraction Overlap (DO, -0.02) and an enhancement in Temporal Understanding (TU, +0.1). This indicates that Vid-LLMs equipped with NumPro maintain robust performance in general video-QA while excelling in precise video temporal grounding (VTG) tasks. This dual capability allows us to harness a powerful Vid-LLM for end-to-end video understanding that can flexibly adapt to both general and temporally nuanced questions within conversational AI systems. Moreover, we examine NumPro on more video-QA benchmarks including MVBench [37] and VideoMME [16], and we show Vid-LLMs enhanced with NumPro achieve robust performance across a variety of downstream tasks. Details can be found in Appendix 9.
@@ 5 Conclusion
Acknowledgement
This work is supported by the National Science Foundation of China (62206048), the Natural Science Foundation of Jiangsu Province (BK20220819), the Fundamental Research Funds for the Central Universities (2242024k30035), and the Southeast University Big Data Computing Center.
@@ 7.2 VTG-Tuned Vid-LLMs
Table 5: The ablation results on the QVHighlights dataset.
@@ 8 Additional Attention Analysis
9 Additional Video Benchmark Results
We conducted experiments on additional video question answering (QA) benchmarks, MVBench [37] and VideoMME [16], as summarized in Table 8. We use 1FPS as the sampling rate, and adopt a design of red color, font size 40, and bottom right positioning for the number prompt. Our results demonstrate that Vid-LLMs enhanced with NumPro achieve robust performance across diverse downstream tasks. Notably, NumPro significantly improves the Vid-LLMs’ generalization capabilities in temporal understanding tasks, such as Scene Transition and Temporal Perception. These findings align with our previous results presented in Table 4 on VideoInstruct [50] in the main paper.
Table 8: Evaluations on two video QA benchmarks: MVBench and VideoMME. The results demonstrate that our NumPro approach enhances Vid-LLMs’ generalization capabilities on downstream tasks involving temporal understanding.
MVBench Video-MME
Scene Transition State Change Overall Temp. Per. Temp. Rea. Overall
$80.0_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+2.5)}}}$ $42.0_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+1.0)}}}$ $51.8_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+0.2)}}}$ $72.7_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+12.7)}}}$ $49.7_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+8.8)}}}$ $63.7_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+0.3)}}}$
@@ 11 Ablation Results on NumPro-FT Designs
Table 9: Comparison between overlaying timestamps with overlaying frame numbers in NumPro design.
Dataset R@0.3 R@0.5 R@0.7 mIoU
Charades-STA $58.4_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-2.3)}}}$ $37.8_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+1.0)}}}$ $16.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+0.7)}}}$ $37.6_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-0.9)}}}$
ActivityNet $31.6_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-12.6)}}}$ $18.1_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-6.3)}}}$ $10.2_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-4.2)}}}$ $23.0_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-8.3)}}}$
12 NumPro with Accurate Timestamps
In the main paper, we choose frame numbers because they serve as the smallest discrete units of a video and can be directly mapped to precise timestamps using the frame sampling rate. In this section, we compare the performance by directly overlaying actual timestamps in videos. However, timestamps (e.g., “10.5s”) may introduce decimals, which can increase parsing complexity for Vid-LLMs. In Tabel 9, we compare minute-level temporal annotations (e.g., “01:10”) with frame numbers annotations sampled with 1FPS. The performance on Charades-STA is comparable, while frame numbers outperformed timestamps on ActivityNet, which includes longer videos and more annotations. These suggest that overlaying numbers with temporal information is an effective strategy for Vid-LLMs in temporal grounding, while frame numbers offer a simpler and more scalable solution.